4. Zinc metal reacts with hydrochloric acid according to the balanced equation Zn(s) + 2 Ha (aq) ZnCl2 (aq) + H2 (g) - → When 0.103 g of Zn(s) is combined with enough HCI to make 50.0 mL of solution in a coffee-cup calorimeter, all of the zinc reacts, raising the temperature of the solution from 22.5°C to 23.7°C. Find 솨lan for this reaction as written. (Use 1.0 g/mL for the density of the solution and 4.18 J/g°C as the specific heat capacity.)
Answer #1

mol of Zn = mass/MW = 0.103/65.39 = 0.001575 mol

Q = m*C*(dT)

Q = 50*4.18*(23.7-22.5)

Q = 250.8 J

HRxn = -Q / n = -250.8/0.001575

HRxn = -159238.095 J/mol

HRxn. =-159.24 kJ/mol